# REAP the Experts: Why Pruning Prevails for One-Shot MoE compression

### v1

@@ 3.3 Empirical evidence for loss of independent control
We analyse the functional subspaces of expert outputs across four diverse state-of-the-art SMoE architectures by recording mean expert activations from 32 samples of 2048 tokens from the c4 dataset (Raffel et al., 2020). By projecting expert activations onto their first two principal components, we visualize how pruning and merging affect the learned representations. See Appendix˜B for additional discussion.
Early vs. late behaviour.
Figures˜1 and A4 demonstrate a striking progression of functional collapse from early to late layers across all architectures. In early layers, the original experts form relatively compact manifolds with moderate spread. After pruning, the surviving experts maintain their positions on the original manifold, preserving its geometric structure with reduced density. In contrast, merging produces a visible contraction toward the manifold’s centre. The contrast becomes dramatic in late layers, where experts are more specialized, and in high granularity architectures with many experts per layer.
@@
Synthesis across architectures.
The consistency of these patterns across architectures with vastly different expert counts (8 to 128), sparsity levels (6.25% to 25% active), and parameter scales (21.9B to 109B) demonstrates that functional collapse under merging is a fundamental property of the operation rather than an artifact of specific implementations. These visualizations reveal that the core issue is not the reduction in the number of experts per se, but rather the qualitative change in the router’s control structure.
@@ 4 Router-weighted Expert Activation Pruning (REAP)
The above analysis demonstrates that the functional output space of a SMoE layer is defined by the coordinated behaviour of the router and experts. An expert’s total contribution to its layer’s output is determined by both its gate-value, $g_{k}(x)$, and the magnitude of its output vector, $\big\|f_{k}(x)\big\|_{2}$. However, naive frequency-based pruning fails to consider these properties. Intuitively, pruning experts which contribute minimally to the layer output minimizes the difference between the original and pruned layer outputs. Let $h(x)$ be the original output and $\bar{h}_{\setminus j}(x)$ be the output after pruning expert $j$ and re-normalizing the remaining router weights. The error induced by pruning expert $j$ is
$\Delta\bar{h}_{\setminus j}(x):=h(x)-\bar{h}_{\setminus j}(x)=\sum_{k}g_{k}(x)f_{k}(x)-\sum_{k\neq j}\frac{g_{k}(x)}{1-g_{j}(x)}f_{k}(x).$ (7)
Re-normalization of the router weights after pruning expert $j$ modulates all other remaining expert outputs, making direct minimization of $\Delta h_{j}$ complex. However, since our goal is to prune unimportant experts, we can reasonably assume their gate-values are small when active $\mathbb{E}_{x\sim\mathcal{X}}[g_{j}(x)]\ll 1$. Under this assumption, the weight re-normalization factor is negligible, i.e., $1-g_{j}(x)\approx 1$, and the error induced by pruning expert $j$ is approximately equal to the expert’s direct contribution to the layer output
$\Delta\bar{h}_{\setminus j}(x)\approx\sum_{k}g_{k}(x)f_{k}(x)-\sum_{k\neq j}g_{k}(x)f_{k}(x))=g_{j}(x)f_{j}(x).$ (8)
To select which experts to prune, we propose a novel saliency criterion, REAP, which approximates an expert’s importance by measuring its direct contribution to the layer’s output magnitude. Specifically, the saliency score, $S_{j}$, is defined as the average of this contribution over tokens for which the expert is active
where $S_{j}$ is the saliency of expert $f_{j}$ and $\mathcal{X}_{j}$ is the set of inputs where $g_{j}(x)\in TopK(\mathbf{g(x))}$.
@@
where $\mathcal{X}_{j}$ is the set of tokens where expert $j$ is active (i.e., $\mathcal{X}_{j}=\{x\mid j\in\text{TopK}(\mathbf{g}(x))\}$). The experts with the minimum saliency score are selected for pruning. REAP is robust to outlier activations and has a direct, intuitive interpretation by quantifying the average magnitude an expert adds to the output vector when it is selected by the router. Pruning experts with the lowest $S_{j}$ removes those with the least impactful contribution.
@@ 5 Experiments
Model
Routed
Experts
Shared
Experts
Top-K
Sparsity
Parameters
Active
Params. (1e9)
First layer
dense
ERNIE-4.5-21B-A3B-PT 64 2 6 87.88% 21.9 3 Yes
Qwen3-30B-A3B 128 0 8 93.75% 30.5 3 No
Mixtral-8x7B-Instruct-v0.1 8 0 2 75.00% 46.7 13 No
GLM-4.5-Air 128 1 8 93.02% 106.9 12 Yes
Llama-4-Scout-17B-16E-Instruct 16 1 1 88.24% 107.8 17 No
Qwen3-Coder-480B-A35B-Instruct-FP8 160 0 8 95.00% 480.2 35 No
Kimi-K2-Instruct-W4A16 (RedHatAI, 2025) 384 1 8 97.66% 1026.4 32 Yes
@@
Compressed SMoE models are evaluated on a suite of benchmarks including MC question answering, code generation, mathematical reasoning, creative writing, and tool calling. See Appendix˜C for details. We implement HC-SMoE and M-SMoE as expert merging baselines. Average linkage criterion is used for HC-SMoE. M-SMoE does not include low-rank compression from the complete MC-SMoE method. Pruning baselines consist of frequency-based pruning and EAN. See Appendix˜D for formal definitions.
@@ 5.1 Results
In Tables˜2 and 2 code generation, creative writing, math reasoning, and MC results are presented for Qwen3-30B and GLM-4.5-Air after calibration with the evol-codealpaca dataset. Table˜3 contains results for large-scale SMoE pruned models on code generation, tool calling, and MC benchmarks. See Table˜A4 and Table˜A5 for detailed MC and code generation results, respectively. Figure˜A5 depicts coding generation and MC accuracy verses model parameters. See Appendix˜E for additional results.
@@
Both merging and pruning are capable of producing accurate compressed SMoE models for MC question answering. HC-SMoE and REAP have a mean decrease in accuracy of approximately 4% and 13% for compression ratios of 25% and 50%, respectively, excluding large-scale SMoEs. REAP achieves first or second rank among all methods, models and compression ratios, suggesting strong consistency regardless of specific model architecture. When calibrated on c4, we find slightly improved accuracies for all compression methods with similar rankings as noted above, see Table˜A6.
@@
Compared to MC, generative benchmarks are more representative of real-world use cases of LLMs. In this setting, pruning emerges as the clearly superior compression method on the generative task benchmarks. Excluding large-scale SMoEs, REAP achieves a mean decrease in accuracy of 2.8% and 8.0% at 25% and 50% compression ratios, respectively, on coding. In comparison, both HC-SMoE and M-SMoE produce mean decreases in accuracy >5% at 25% compression and >20% at 50% compression. Notably, REAP maintains significantly higher accuracy at 50% compression than other pruning methods. On creative writing, REAP and EAN are near-lossless at 25% compression with REAP offering improved quality at 50% compression. Merging methods are less consistent across various model architectures and compression ratios. For example, M-SMoE is the best method for Qwen3-30B at 50% compression, but the worst on GLM-4.5-Air. REAP attains the best mathematical reasoning results with a remarkable mean decrease in accuracy of just 1.1% at 50% compression. HC-SMoE and M-SMoE offer high accuracy at 25% compression but are significantly less accurate than pruning at 50% compression.
@@
Qwen3-30B-A3B Baseline 0.859 0.302 0.581 0.811 0.903 0.872 0.887 0.721
25% Merging M-SMoE 0.822 0.293 0.558 0.805 0.901 0.872 0.886 0.558
HC-SMoE 0.800 0.258 0.529 0.497 0.864 0.834 0.849 0.674
Pruning Frequency 0.849 0.302 0.576 0.807 0.905 0.864 0.885 0.600
EAN 0.840 0.311 0.576 0.811 0.895 0.866 0.881 0.603
REAP 0.843 0.308 0.575 0.804 0.892 0.864 0.878 0.669
50% Merging M-SMoE 0.621 0.205 0.413 0.725 0.824 0.838 0.831 0.451
HC-SMoE 0.574 0.185 0.379 0.008 0.760 0.696 0.728 0.542
Pruning Frequency 0.704 0.236 0.470 0.677 0.882 0.860 0.871 0.483
EAN 0.798 0.306 0.552 0.702 0.886 0.842 0.864 0.493
REAP 0.821 0.293 0.557 0.718 0.878 0.872 0.875 0.518
GLM-4.5-Air Baseline 0.820 0.374 0.597 0.839 0.846 0.918 0.882 0.747
25% Merging M-SMoE 0.781 0.330 0.555 0.781 0.848 0.880 0.864 0.596
HC-SMoE 0.793 0.363 0.578 0.788 0.842 0.908 0.875 0.704
Pruning Frequency 0.805 0.341 0.573 0.793 0.832 0.908 0.870 0.648
EAN 0.821 0.374 0.597 0.824 0.839 0.908 0.874 0.637
REAP 0.794 0.390 0.592 0.831 0.835 0.926 0.880 0.678
50% Merging M-SMoE 0.493 0.099 0.296 0.391 0.465 0.466 0.465 0.444
HC-SMoE 0.662 0.220 0.441 0.593 0.667 0.732 0.700 0.564
Pruning Frequency 0.546 0.104 0.325 0.604 0.615 0.612 0.613 0.521
EAN 0.773 0.253 0.513 0.702 0.781 0.838 0.809 0.511
REAP 0.755 0.352 0.553 0.754 0.820 0.926 0.873 0.559
@@
Qwen3-Coder-
480B-A35B-
Instruct-FP8
Baseline 0.889 0.431 0.660 0.540 0.866 0.825 0.380 0.690 0.750
25% Frequency 0.792 0.296 0.544 0.378 0.844 0.763 0.355 0.654 0.606
EAN 0.876 0.419 0.647 0.534 0.831 0.813 0.384 0.676 0.702
REAP 0.884 0.416 0.650 0.540 0.878 0.823 0.392 0.698 0.748
50% Frequency 0.011 0.012 0.011 0.000 0.200 0.392 0.000 0.197 0.506
EAN 0.831 0.382 0.607 0.536 0.822 0.774 0.383 0.659 0.591
REAP 0.873 0.415 0.644 0.522 0.849 0.801 0.371 0.674 0.692
Kimi-K2-
Instruct-
W4A16
Baseline 0.883 0.434 0.659 0.554 0.840 0.802 0.355 0.666 0.780
25% Frequency 0.524 0.082 0.303 0.000 0.644 0.603 0.045 0.431 0.604
EAN 0.831 0.379 0.605 0.562 0.819 0.802 0.335 0.652 0.703
REAP 0.889 0.440 0.664 0.580 0.842 0.801 0.263 0.635 0.773
50% Frequency 0.124 0.000 0.062 0.000 0.255 0.397 0.003 0.218 0.439
EAN 0.772 0.253 0.513 0.576 0.778 0.767 0.173 0.573 0.587
REAP 0.863 0.429 0.646 0.576 0.785 0.743 0.164 0.564 0.643
@@
To asses whether pruning remains viable at scale, we prune Qwen3-Coder-480B and Kimi-K2-Instruct. On MC questions, REAP outperforms other pruning methods. On non-agentic coding tasks, REAP achieves near-lossless accuracy with a 0.20% and 1.4% mean decrease in accuracy compared to baseline at 25% and 50%, respectively, outperforming EAN and frequency-based pruning, particularly at 50% compression. On the challenging SWE-Bench task, both REAP and EAN maintain high accuracy at 25% and 50% compression, with some scores slightly exceeding the baseline. On tool use, EAN and REAP are comparable, with REAP slightly outperforming at 50% compression with a mean decrease in accuracy of 5.9% versus 6.2% for EAN. Frequency-based pruning suffers from a sharp degradation in quality at 50% compression, highlighting the importance of pruning saliency criteria which consider expert activations. Scaling the pruning methods is relatively trivial. Unlike HC-SMoE, calibration for pruning does not require recording activations from every expert for every token, facilitating efficient calibration. Further, pruning can be easily applied to quantized models without any additional steps required to reconcile block scales or re-quantize following compression.
@@
Figure 3: (3(a)) & (3(b)) N-Gram diversity and cross-perplexity of compressed Qwen3-30B-A3B models at 50% compression, respectively. (3(c)) JSD of compressed and baseline model logits vs. completion token position for Qwen3-30B-A3B at 50% compression. Initially, all compressed models share close alignment with the baseline model. However, as the completion token position increases the merged models diverge from the baseline more rapidly than the REAP pruned model. (3(d)) The mean relative L2-distance and singular-vector alignment between Qwen3-30B expert weights at 50% compression. Expert merging is more challenging than model merging due to large distances between experts in weight space and low singular-vector alignment.
Quantifying merged MoE generation quality.
While merged expert SMoEs offer reasonable quality for discriminative tasks such as MC question and answering, they fail to remain competitive on generative tasks. To help explain the performance gap of merged models between discriminative and generative tasks, we perform an analysis of the compressed model outputs and compare with REAP pruned models. We prompt 50% compressed Qwen3-30B models with 100 questions randomly sampled from the evol-codealpaca dataset and record their outputs. In Figure˜3(a), we measure the N-gram diversity and find that the merged models have significantly lower diversity across all N-gram sizes measured. In contrast, the REAP pruned model remains similar to the base model, albeit slightly less diverse. In Figure˜3(b), we measure the perplexity of the text generated by the compressed models with the original baseline model. The text generated by the merged models has both a higher mean and higher variance than the pruned model generations, suggesting that the REAP pruned model outputs are more closely aligned to the original model. The alignment between the baseline and REAP pruned SMoEs is further supported by Figure˜3(c), which plots the JSD of the compressed and baseline logits vs. output token position. The merged model logits diverge from the baseline more rapidly than the pruned model.
@@
Model merging has been widely adopted to facilitate LLM fine-tuning. Why does expert merging miss the mark? In addition to the loss of the router’s input-dependent modulation of experts explored in Section˜3, we argue that the non-local nature of expert merging and high cardinality of expert clusters pose significant unresolved challenges.
In Figure˜3(d), we plot the mean relative L2-distance between experts clustered by HC-SMoE or M-SMoE and compare with the distance between expert weights from the pretrained to Instruct Fine-Tuned (IFT) checkpoints. We find that the distance between clustered experts within the same layer greatly exceeds that of experts in the IFT checkpoint after fine-tuning. Ito et al. (2024) found that weight matching permutations improved alignment of parameters’ singular vectors. Following their approach, we decompose expert weights with Singular Value Decomposition (SVD) and plot the singular-vector alignment in Figure˜3(d). Even after applying weight matching permutations, the M-SMoE expert clusters remain far apart both in weight space and singular-vector alignment. The relatively poorly aligned experts highlight the considerable challenge of coherently merging their parameters.
When merging works well, it’s more closely related to pruning than one might expect. In Figure˜6(a), we depict the frequency of singleton clusters — clusters containing a single expert — for both HC-SMoE and M-SMoE. A singleton cluster is directly analogous to an expert that remains after pruning. We find that HC-SMoE in particular has a high prevalence of singleton clusters, leaving important experts unadulterated and compressing the rest into a few mega-clusters containing tens of experts. This is particularly true of the high granularity models which contain more experts per layer. We hypothesize that the cardinality of these mega-clusters poses a challenge for existing merging algorithms and test this intuition in Figure˜6(b). Unfortunately, even modest restrictions of the maximum cluster size to 32 — half the number of experts to compress — results in large decreases in model quality on coding tasks.
@@
In Figure˜A7, we plot the code generation accuracy of the various compression methods and models when calibrated on either c4 or evol-codealpaca. The difference is stark, c4 calibration results in a collapse in accuracy, with several compressed model instances failing to produce coherent outputs, resulting in 0% accuracy. In Figure˜A8, we compare the accuracy of compressed Qwen3-30B models calibrated with either domain-specific data or the combined calibration data across all generative tasks. The domain-specific calibrated models achieve significantly higher accuracy, especially at 50% compression.

### v2

@@ 3.3 Empirical evidence for loss of independent control
We analyse the functional expert output manifolds across four diverse state-of-the-art SMoE architectures by recording mean expert activations from 32 samples of 2048 tokens from the C4 dataset (Raffel et al., 2020).
Functional subspace collapse.
By projecting expert activations onto their first two principal components, we visualize how pruning and merging affect the learned representations. Figures˜1, 5(a) and 5(b) demonstrate a striking progression of functional subspace collapse from early to late layers in high-granularity architectures such as Qwen3 and ERNIE-4.5. In early layers, the original experts form relatively compact manifolds with moderate spread. After pruning, the surviving experts maintain their positions on the original manifold, preserving its geometric structure with reduced density. In contrast, merging produces a visible contraction toward the manifold’s centre. The contrast becomes dramatic in late layers, where experts are more specialized.
@@
Functional manifold distortion.
While collapse is less apparent in low-granularity models, the introduction of novel functions due to merging distorts the topology of the original expert manifold to a greater degree than pruning. To quantitatively measure this phenomenon, we measure the $1$-Wasserstein distance (Earth Mover’s distance) between the original and compressed expert output manifolds, see Section˜B.2 for details. As depicted in Figure˜2, the merged outputs consistently exhibit a higher transport cost from the original manifold.
Manifold geometry preservation.
Across all models and layers, we observe that pruning preserves the topology of the functional manifold while merging fundamentally alters it. The preservation of manifold geometry under pruning reflects the mathematical structure of the operation: the pruned expert class is a coordinate subspace of the original, with the router maintaining independent control over each surviving expert.
In contrast, the subspace collapse observed in merged highly-granular SMoEs visualizes the loss of independent control. When gates $g_{i}$ and $g_{j}$ are tied by their sum $(g_{i}+g_{j})$, the router can no longer independently modulate the two underlying functions, forcing the model to approximate the dynamic mixture $r(x)f_{i}(x)+(1-r(x))f_{j}(x)$ with a static merged expert $\tilde{f}$.
With low-granularity SMoEs, such as Llama-4-Scout and Mixtral, functional subspace collapse due to expert merging is less apparent, see Figures˜5(c), 5(d), 5(e) and 5(f). With few experts per layer and active experts per token, these architectures have less variable routing and higher gate-values, which better preserves the variance of the original manifold. However, the introduction of novel functions by merging introduces greater manifold distortion than the substitution error associated with pruning. These observations reveal that the core issue is not the reduction in the number of experts per se, but rather the qualitative change in the router’s control structure and the introduction of novel functionality. See Appendix˜B for additional discussion.
@@
(a) Qwen3-30B
(b) ERNIE-4.5-21B
(c) Mixtral-8x7B
(d) Llama-4-Scout
Figure 2: 11-Wasserstein distance between the compressed and original expert output manifolds measured in normalized angular distance. Expert merging introduces novel functions which distort the manifold.
@@ 4 Router-weighted Expert Activation Pruning (REAP)
The motivation in Section˜3 demonstrates that the functional output space of a SMoE layer is defined by the coordinated behaviour of the router and experts. As established in Equation˜8, the magnitude of the substitution error incurred by promoting expert $i$ in lieu of pruned expert $j$ is upper bounded by $g_{j}(x)(\|f_{j}(x)\|+\|f_{i}(x)\|)$. Naive frequency-based pruning considers neither the coordination between router and expert ($g_{j}(x)$) nor the functional properties of the pruned expert ($\|f_{j}(x)\|)$, effectively assuming that all active experts contribute equally to the output. By ignoring these terms, frequency-based methods fail to minimize the error bound derived above.
Since the identity of the promoted expert $i$ (and thus $\|f_{i}(x)\|$) varies across tokens, directly minimizing the pruned expert’s impact $g_{j}(x)\|f_{j}(x)\|$ is an effective heuristic to minimize the total error. This strategy targets the known components of the error bound ($g_{j}\|f_{j}\|$) while simultaneously shrinking the scaling coefficient ($g_{j}$) of the unknown component ($\|f_{i}\|$). Intuitively, this identifies experts which contribute minimally to the layer output, yielding the minimal difference between the original and pruned layer outputs in expectation. To select which experts to prune, we propose a novel saliency criterion, REAP. Specifically, the saliency score, $S_{j}$, is defined as the average of the expert’s weighted magnitude over tokens for which it is active:
@@
where $\mathcal{X}_{j}$ is the set of tokens where expert $j$ is active (i.e., $\mathcal{X}_{j}=\{x\mid j\in\mathcal{T}(x)\}$). Crucially, calculating this average conditionally over $\mathcal{X}_{j}$ rather than globally decouples the expert’s functional impact from its frequency of activation. A global average may be dominated by usage frequency and risks pruning specialist experts which are rarely activated but contribute significantly to the layer output when selected. By pruning experts with the lowest $S_{j}$, REAP targets those that provide a weak functional contribution even when specifically requested by the router, thereby minimizing the substitution error bound for every active token.
@@ 5 Experiments
Model Routed Experts Shared Experts Top-K Sparsity Parameters (1e9) Active Params. (1e9) First layer dense Expert Granularity
ERNIE-4.5-21B-A3B-PT 64 2 6 87.88% 21.9 3 Yes High
Qwen3-30B-A3B 128 0 8 93.75% 30.5 3 No High
Mixtral-8x7B-Instruct-v0.1 8 0 2 75.00% 46.7 13 No Low
GLM-4.5-Air 128 1 8 93.02% 106.9 12 Yes High
Llama-4-Scout-17B-16E-Instruct 16 1 1 88.24% 107.8 17 No Low
Qwen3-Coder-480B-A35B-Instruct-FP8 160 0 8 95.00% 480.2 35 No High
Kimi-K2-Instruct-W4A16 (RedHatAI, 2025) 384 1 8 97.66% 1026.4 32 Yes High
@@
Compressed SMoE models are evaluated on a suite of benchmarks including MC question answering, code generation, mathematical reasoning, creative writing, and tool calling. See Appendix˜C for details. We implement HC-SMoE and M-SMoE as expert merging baselines. Average linkage criterion is used for HC-SMoE. M-SMoE does not include low-rank compression from the complete MC-SMoE method. Pruning baselines consist of frequency-based pruning and EAN and experts with the lowest saliency scores according to each method’s criterion are pruned. See Appendix˜D for formal definitions.
@@ 5.1 Results
In Tables˜2 and 3 code generation, creative writing, math reasoning, and MC results are presented for Qwen3-30B and GLM-4.5-Air after calibration with domain-specific datasets. Table˜3 contains results for large-scale SMoE pruned models on code generation, tool calling, and MC benchmarks. See Table˜A6 and Table˜A7 for detailed MC and code generation results, respectively. Figure˜A6 depicts coding generation and MC accuracy versus model parameters. See Appendix˜F for additional results.
@@
Both merging and pruning are capable of producing accurate compressed SMoE models for MC question answering. HC-SMoE and REAP have a mean decrease in accuracy of approximately 4% and 13% for compression ratios of 25% and 50%, respectively, excluding large-scale SMoEs. REAP achieves first or second rank among all methods, models and compression ratios, suggesting strong consistency regardless of specific model architecture. When calibrated on C4, we find slightly improved accuracies for all compression methods with similar rankings as noted above, see Table˜A8.
@@
Compared to MC, generative benchmarks are more representative of real-world use cases of LLMs. In this setting, pruning emerges as the clearly superior compression method on the generative task benchmarks. Excluding large-scale SMoEs, REAP achieves a mean decrease in accuracy of 1.9% and 6.9% at 25% and 50% compression ratios, respectively, on coding. In comparison, both HC-SMoE and M-SMoE produce mean decreases in accuracy >5% at 25% compression and >20% at 50% compression. Notably, REAP maintains significantly higher accuracy at 50% compression than other pruning methods. M-SMoE achieves significantly better code generation accuracy on low-granularity SMoE architectures.
On creative writing, REAP and EAN are near-lossless at 25% compression with REAP offering improved quality at 50% compression. Merging methods are less consistent across various model architectures and compression ratios. For example, M-SMoE is the best method for Qwen3-30B at 50% compression, but the worst on GLM-4.5-Air. REAP attains the best mathematical reasoning results with a remarkable mean decrease in accuracy of just 0.1% at 25% compression. In comparison, HC-SMoE and M-SMoE offer high accuracy at 25% compression but are significantly less accurate than pruning at 50% compression.
@@
Qwen3-30B-A3B Baseline 0.814 0.302 0.558 0.811 0.903 0.872 0.887 0.721
25% Merging M-SMoE 0.781 0.293 0.537 0.805 0.901 0.872 0.886 0.558
HC-SMoE 0.752 0.258 0.505 0.497 0.864 0.834 0.849 0.674
Pruning Frequency 0.805 0.302 0.553 0.807 0.910 0.865 0.888 0.600
EAN 0.797 0.311 0.554 0.811 0.904 0.879 0.892 0.603
REAP 0.797 0.304 0.551 0.804 0.896 0.881 0.888 0.665
50% Merging M-SMoE 0.590 0.205 0.397 0.725 0.824 0.838 0.831 0.451
HC-SMoE 0.543 0.185 0.364 0.008 0.760 0.696 0.728 0.542
Pruning Frequency 0.668 0.236 0.452 0.677 0.871 0.860 0.865 0.483
EAN 0.753 0.306 0.530 0.702 0.874 0.855 0.864 0.493
REAP 0.780 0.302 0.541 0.718 0.877 0.838 0.857 0.503
GLM-4.5-Air Baseline 0.786 0.374 0.580 0.839 0.846 0.918 0.882 0.747
25% Merging M-SMoE 0.726 0.330 0.528 0.781 0.848 0.880 0.864 0.596
HC-SMoE 0.737 0.363 0.550 0.788 0.842 0.908 0.875 0.704
Pruning Frequency 0.759 0.341 0.550 0.793 0.832 0.908 0.870 0.648
EAN 0.768 0.374 0.571 0.824 0.839 0.908 0.874 0.637
REAP 0.759 0.412 0.585 0.831 0.839 0.920 0.879 0.674
50% Merging M-SMoE 0.468 0.099 0.284 0.391 0.465 0.466 0.465 0.444
HC-SMoE 0.618 0.220 0.419 0.593 0.667 0.732 0.700 0.564
Pruning Frequency 0.511 0.104 0.308 0.604 0.615 0.612 0.613 0.521
EAN 0.721 0.253 0.487 0.702 0.781 0.838 0.809 0.511
REAP 0.679 0.352 0.515 0.754 0.815 0.900 0.857 0.554
@@
Qwen3-Coder- 480B-A35B- Instruct-FP8 Baseline 0.841 0.431 0.636 0.540 0.866 0.825 0.380 0.690 0.750
25% Frequency 0.737 0.296 0.516 0.378 0.844 0.763 0.355 0.654 0.606
EAN 0.827 0.419 0.623 0.534 0.831 0.813 0.384 0.676 0.702
REAP 0.831 0.416 0.624 0.540 0.878 0.823 0.392 0.698 0.748
50% Frequency 0.007 0.012 0.010 0.000 0.200 0.392 0.000 0.197 0.506
EAN 0.777 0.382 0.580 0.536 0.822 0.774 0.383 0.659 0.591
REAP 0.822 0.415 0.619 0.522 0.849 0.801 0.371 0.674 0.692
Kimi-K2- Instruct- W4A16 Baseline 0.828 0.434 0.631 0.554 0.840 0.802 0.355 0.666 0.780
25% Frequency 0.486 0.082 0.284 0.000 0.644 0.603 0.045 0.431 0.604
EAN 0.779 0.379 0.579 0.562 0.819 0.802 0.335 0.652 0.703
REAP 0.840 0.440 0.640 0.580 0.842 0.801 0.263 0.635 0.773
50% Frequency 0.112 0.000 0.056 0.000 0.255 0.397 0.003 0.218 0.439
EAN 0.722 0.253 0.487 0.576 0.778 0.767 0.173 0.573 0.587
REAP 0.819 0.429 0.624 0.576 0.785 0.743 0.164 0.564 0.643
@@
To assess whether pruning remains viable at scale, we prune Qwen3-Coder-480B and Kimi-K2-Instruct. On MC questions, REAP outperforms other pruning methods. On non-agentic coding tasks, REAP achieves near-lossless accuracy with a 0.16% and 1.2% mean decrease in accuracy compared to baseline at 25% and 50%, respectively, outperforming EAN and frequency-based pruning, particularly at 50% compression. On the challenging SWE-Bench task, both REAP and EAN maintain high accuracy at 25% and 50% compression, with some scores slightly exceeding the baseline. On tool use, EAN and REAP are comparable, with REAP slightly outperforming at 50% compression with a mean decrease in accuracy of 5.9% versus 6.2% for EAN. Frequency-based pruning suffers from a sharp degradation in quality at 50% compression, highlighting the importance of pruning saliency criteria which consider expert activations. Scaling the pruning methods is relatively trivial. Unlike HC-SMoE, calibration for pruning does not require recording activations from every expert for every token, facilitating efficient calibration. Further, pruning can be easily applied to quantized models without any additional steps required to reconcile block scales or re-quantize following compression. See Appendix˜E for further discussion.
@@
Figure 4: (4(a)) & (4(b)) N-Gram diversity and cross-perplexity of compressed Qwen3-30B-A3B models at 50% compression, respectively. (4(c)) JSD of compressed and baseline model logits vs. completion token position for Qwen3-30B-A3B at 50% compression. Initially, all compressed models share close alignment with the baseline model. However, as the completion token position increases the merged models diverge from the baseline more rapidly than the REAP pruned model. (4(d)) The mean relative L2-distance and singular-vector alignment between Qwen3-30B expert weights at 50% compression. Expert merging is more challenging than model merging due to large distances between experts in weight space and low singular-vector alignment.
Quantifying merged SMoE generation quality.
While merged expert SMoEs offer reasonable quality for discriminative tasks such as MC question and answering, they fail to remain competitive on generative tasks. To help explain the performance gap of merged models between discriminative and generative tasks, we perform an analysis of the compressed model outputs and compare with REAP pruned models. We prompt 50% compressed Qwen3-30B models with 100 questions randomly sampled from the evol-codealpaca dataset and record their outputs. In Figure˜4(a), we measure the N-gram diversity and find that the merged models have significantly lower diversity across all N-gram sizes measured. In contrast, the REAP pruned model remains similar to the base model, albeit slightly less diverse. In Figure˜4(b), we measure the perplexity of the text generated by the compressed models with the original baseline model. The text generated by the merged models has both a higher mean and higher variance than the pruned model generations, suggesting that the REAP pruned model outputs are more closely aligned to the original model. The alignment between the baseline and REAP pruned SMoEs is further supported by Figure˜4(c), which plots the JSD of the compressed and baseline logits vs. output token position. The merged model logits diverge from the baseline more rapidly than the pruned model.
@@
Model merging has been widely adopted to facilitate LLM fine-tuning. Why does expert merging miss the mark? In addition to the loss of the router’s input-dependent modulation of experts explored in Section˜3, we argue that the non-local nature of expert merging and high cardinality of expert clusters pose significant unresolved challenges. In Figure˜4(d), we plot the mean relative L2-distance between experts clustered by HC-SMoE or M-SMoE and compare with the distance between expert weights from the pretrained to Instruct Fine-Tuned (IFT) checkpoints. We find that the distance between clustered experts within the same layer greatly exceeds that of experts in the IFT checkpoint after fine-tuning. Ito et al. (2024) found that weight matching permutations improved alignment of parameters’ singular vectors. Following their approach, we decompose expert weights with Singular Value Decomposition (SVD) and plot the singular-vector alignment in Figure˜4(d). Even after applying weight matching permutations, the M-SMoE expert clusters remain far apart both in weight space and singular-vector alignment. The relatively poorly aligned experts highlight the considerable challenge of coherently merging their parameters.
When merging works well, it’s more closely related to pruning than one might expect. In Figure˜7(a), we depict the frequency of singleton clusters — clusters containing a single expert — for both HC-SMoE and M-SMoE. A singleton cluster is directly analogous to an expert that remains after pruning. We find that HC-SMoE in particular has a high prevalence of singleton clusters, leaving important experts unadulterated and compressing the rest into a few mega-clusters containing tens of experts. This is particularly true of the high granularity models which contain more experts per layer. We hypothesize that the cardinality of these mega-clusters poses a challenge for existing merging algorithms and test this intuition in Figure˜7(b). Unfortunately, even modest restrictions of the maximum cluster size to 32 — half the number of experts to compress — results in large decreases in model quality on coding tasks.
@@
In Figure˜A8, we plot the code generation accuracy of the various compression methods and models when calibrated on either C4 or evol-codealpaca. The difference is stark, C4 calibration results in a collapse in accuracy, with several compressed model instances failing to produce coherent outputs, resulting in 0% accuracy. In Figure˜A9, we compare the accuracy of compressed Qwen3-30B models calibrated with either domain-specific data or the combined calibration data across all generative tasks. While domain-specific calibration results in higher accuracies, REAP best preserves accuracy compared to other compression methods in the combined data calibration setting.